add 2 1/2, 3/4, and 1/3 together And please show how it is done

Answer by checkley77(12844)

add 2 1/2, 3/4, and 1/3
2 1/2+3/4+1/3
First we change the mixed fraction 2 1/2 to an improper fraction.
(2*2+1)/2=(4+1)/2=5/2
5/2+3/4+1/3
Now we find a common denominator so we can combine these fractions.
3*4=12 is the LCD.
(5*6+3*3+1*4)/12
Now we add the numerators.
(30+9+4)/12
43/12 ans.